OpenStudy (anonymous):

A group of college students is volunteering for Homes for the Community during their spring break. They are putting the finishing touches on a house they built. Working alone, Wade can paint a certain room in 4 hours. Rhonda can paint the same room in 6 hours. How long will it take them working together to paint the room? Round your answer to the nearest hundredth if necessary. A. 2.4 hours B. 0.1 hours C. 5 hours D. 12 hours @Cpt.Sparz @LadyKatsaTheGraceling

OpenStudy (anonymous):

I THINK D.12 HOURS

OpenStudy (anonymous):

T = 12/7... and that will be the answer

OpenStudy (anonymous):

Wait don't follow this I was looking at it wrongly...!!!

OpenStudy (anonymous):

? i got "1 r 5" using long division and when i evaluated it i got "1.714285714286"

OpenStudy (anonymous):

Yea...I did the math wrong!

OpenStudy (anonymous):

lol, take ur time

OpenStudy (anonymous):

The answer is C 5 hours

OpenStudy (anonymous):

how? can u show the work, u don't have to if u don't want

OpenStudy (anonymous):

t/6 + t/4 = hour 6t + 4t = 24t or 10t = 24 so t = 24/10

OpenStudy (anonymous):

2.4 so the answer is A

OpenStudy (anonymous):

wow just realized I put is instead of isnt in front 5 hours! sorry man

OpenStudy (anonymous):

oh, i thought it was D because 6*4=24/2=12

OpenStudy (anonymous):

Yea I did that at first but I was like that doesn't seem right so I checked out other sources and found how to do it

OpenStudy (anonymous):

so it is A. right?

OpenStudy (anonymous):

Yeppers! :3

OpenStudy (anonymous):

k thanks man.

OpenStudy (anonymous):

24/10 = 2.4

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

I say c. 5 cause if you find the average of both wade and rhonda and add them together I got 5.

OpenStudy (anonymous):

wade's average was one wall per hour and rhonda's was one wal per 1.5 combine their averages and divide by two and you get 1.25. this means that together they can finish one wall in one hour and 25min. so multiply 1.25 x 4 = 5. thus proving my answer!

OpenStudy (anonymous):

im confused!! hold on